yeetantor:

If a piece of ribbon is 6 and 1 over 4 feet long how long is it in inches

yeetantor:

its actually math btw

Angelwolf:

So first off we know 1 ft = 12 inches so 6 ft = 6 * 12 = 72 inches 1/4 of a ft = 1/4(12) = 3 inches so your answer is 75 inches